Monday, September 30, 2013

Night Log 2013-09-29

SA: Amanda
SO: Veronica
Others: Janus, Hamish, HRS team, David


. Dark night with good conditions.

. Data taken for the following programs:
2013-1-RSA_HET_UKSC_OTH-001 (P0 BVIT)
HRS commissioning

Sunday, September 29, 2013

Night Log 2013-09-28

SA: Amanda
SO: Veronica
Others: HRS team (Ray, David, Jurgen, Lisa, Luke), David, Janus, Chris, Keith, Hamish, Francois


. First light for HRS!! All data taken tonight was commissioning our spiffy new instrument. See the blog for details and photos.

. Dark night with excellent seeing (subarcsec for a few hours). We closed ~02:00 due to high humidity.

Saturday, September 28, 2013

HRS First Light!!!

It's Saturday morning so it's Tracker Time!  It's a testament to Jürgen's endless enthusiasm that he somehow managed to get into his safety harness, despite almost every bit of it being thoroughly twisted.  We did get him straightened out eventually though ;)

Up top, Eben & Jono finished getting the fibre cable into the rho wrap.

Ray made sure the fibres emerging from the break-out box were as comfortable & safe as possible, but we'll have to get Tech Ops to make us a cover of some sort to protect this awfully vulnerable part of the instrument.

During lunch, Jono & Johan removed the hosepipe & attached all the clips & brackets that support the fibre cable.

After lunch, Janus installed the FIF in the payload.

So then the fibres could be fed through from the top of the rho stage, into the FIF bay in the payload's Rotating Structure.

As anticipated, this was not a trivial task & some tweaks were needed before we could get all the ferrules safely bolted into position.

Eben made a cool holster (the blue tube on the left) to contain & protect the spare fibres.  This would also serve as a parking bay for the actual (non-spare) fibres if for whatever reason we had to remove the FIF.  We checked that all the fibres had enough slack to comfortably reach all the extreme FIF positions & everything looks Good!

The team came back up after a quick dinner, fully expecting the humidity to spike at any time & shut things down.

Aligning the primary was a real mission as we'd had the aircons off for most of the afternoon while working up on the tracker.

We had to settle for a coarse alignment in the end, but at least the weather held up & we could set about initialising & configuring the mechanisms for our first on-sky exposure.

We wanted a bright target & hoped for something auspicious like Alpha Cen, but nothing of the sort was within SALT's visibility annulus at the time.  So we pointed to HD190007, a V = 7.5 mag K4 dwarf that's a high-precision radial velocity standard.

No shortage of tension in the control room...

Janus sent the Low Res object fibre to fetch the star & we set the exposure meter running.  Counts!!  But are they really coming from the star?  Move the FIF away slightly - see the counts drop - move back - see them return: Yes :D  We have starlight going down a HRS fibre!

Sorry Ray, but it really is time for the PI to wear The Astronomer's Hat!

A 10 second exposure, a 12 second readout & we have our first blue échellogram - Her Royal Spectrographness Is Alive!! :D

& the red channel's working too - here's a 2 minute High Res exposure of HD199961, a V = 4.7 mag G6 giant.

SALT Astro & Tech Ops Managers are smiling: "I'm happy if You're happy!" says Chris :)

Just before snatching my camera away.

With that, Ray disappeared to the fridge to fetch a bottle of bubbly & was again required to don The Hat.  

Pity Eddy & Ockert & various others aren't here for this, but Cheers to all!  & especially to Eben & the guys that worked all day (through lunch & dinner on a Saturday) to allow us to go on-sky tonight while the weather's good, before Jürgen & David leave.  Also to Amanda & Veronica for so graciously letting us crash your observing run!

We have another treasure to add to the collection on top of the SALT fridge...

The weather continued to amaze as the humidity levelled off & the seeing steadily improved - we don't often see 0.6 arcsecond seeing up here!

So inspiring was all this that next thing we found Hamish grilling Francois about the finer details of how the CCAS instrument works to align the primary mirror.

Through the rest of the night we observed a few other objects in the different modes & tried various calibration frames, including the ThAr arc lamp in the SALT Calsys.  We have heaps more to do, but for now we're finding our feet & figuring out what's what...

Along the way, the control room crowd slowly thinned out.

& we began dropping like flies through the wee hours...

Jürgen might well be sleep-walking in this one?  David's focused at infinity while Ray (of course), is actually thinking.

Shortly after pointing out that the control room sorely needs a hammock, Luke checked out.

Finally even Ray went down...  That's when you know it's time to stop!

It's been a very long & busy 21 hour day for us - we're off to bed...  Sleep tight HRS :)

Night Log 2013-09-27

SA: Amanda
SO: Veronica
Others: Janus, Pierre, David


. No data tonight. Bad weather and telescope offline for HRS fiber work.

Friday, September 27, 2013

Up Go The Fibres!!

The first task for today was to set up a toy "star" to be imaged onto the FIF's acquisition camera.  This was done using a blue LED to feed a pinhole & a diffuser placed in front of a microscope objective.  The star was defocused to produce a large spot with which the stage positions for the various fibres could be optimised.

So then it was time for Jürgen to start attaching the object & sky fibres to the FIF.

The brass ferrules for the object (left) & sky (right) fibres for the High Stability (HS) mode can be seen just below the CCD chip (near the centre of the photo).  This is at the FIF's maximum separation for a mode pair, 29 mm.

Here's the HS pair again, but at the minimum separation of 4 mm.

As above, seen up close...

Installing the fibres was rather a fiddly process, but it's guaranteed to be even fiddlier when doing this up on the tracker!

This shows all the fibres in position, with the FIF jaws at their minimum separation.  From top to bottom the four modes are High Stability (HS), High Res (HR), Medium Res (MR) & Low Res (LR).

This is the view from behind.

With the fibres in place, Janus could monitor the counts on the exposure meter to judge how well the spot's aligned to the fibre being tested. 

Then he could adjust the FIF stages to determine the best position for each fibre & save that encoder reading.  So now the FIF will offset reliably to position the object on the required fibre.

With that final task completed, the fibres could be taken out again & bundled up for their scary ride up to the tracker.  The fibre ends & the break-out box had to be supported together, but some flexibility would be necessary so a bendy piece of board was used & then the whole lot was wrapped in a thin layer of foam.  The fibre ends were also enclosed in plastic tubes for extra protection.

Even IT was willing to get involved in this nerve-wracking process!

One should never look down the barrel...

The team underneath the primary took charge of hauling the precious fibre bundle up through the pintle bearing.

& then carefully feeding it along to the edge of the structure.

Where they could pass the end up to Eben & Johan on the cherry-picker.

Good to see HRS's Papa Bear looking happy at this stage :)

The 50 m long green mamba's now been replaced by a black one...

Apparently this was all pretty amazing to David ;)

This is a great spot to view the primary mirror from.

Having negotiated the loop around the structure node, the aerial guys could proceed up to the tracker.

Safely up on the tracker platform!  Whew :)

Now for the real challenge: getting this whole lot through the rho wrap!  Not a simple matter since only the sides of the cable wrap can be opened up - the steel ends can't be dismantled so everything has to squeeze through the front & back doors...

All the other cables & pipes had to be man-handled out of the way to make space to push the fibres & box through.  

Fortunately the contraption only had to go up through the entrance of the cable wrap & then out the side.  Then the cable could be pushed through & looped back down to the exit end of the rho wrap.

Here it could again come in through the side & then exit through the steel frame at the end of the cable wrap.  

The delicate bits could then be parked up on the rho stage while the rest of the cable was fed in through the side of the cable wrap.  This was a slow process as the plastic slats on the side of the cable wrap had to be removed (before) & replaced (after) pushing the cable into place.  We'll finish this up tomorrow...

Thanks to Janus for contributing various cool photos from today's action, & to Jürgen for this one of the Milky Way & Venus taken the other night.  Extra neat is the Zodiacal Light visible as a bright wedge extending up from the bottom of the image.  This is light scattered off of dust in the plane of our Solar System.  It's particularly obvious at the moment as we're about a week past the equinox so the Zodiac is steeply angled with respect to the horizon.

If all goes well, we should complete the installation of the fibres & the FIF tomorrow.  So if the weather clears up, we'll go on-sky tomorrow night!